Determine whether the given ordered pair is a solution of the given equation.
(x - 5)^2 + (y + 7)^2 = -16; (2, -2)

Answer by checkley77(12844)

(x - 5)^2 + (y + 7)^2 = -16; (2, -2)
(2-5)^2+(-2+7)^2=-16
-3^2+5^2=-16
9+25=-16
34 does not = -16 therefore (2,-2) is not a solution.